OpenStudy (anonymous):

x+=3y=18 , 3x+9y=12 . parallel,perpendicular, coeficient, or none

OpenStudy (anonymous):

im assuming its x+3y...

OpenStudy (anonymous):

lol sorry im having typing errors tonight yeah your correct

OpenStudy (anonymous):

x+3y=18 3x+9y=12 y=-x/3+6 y=-x/3+4/3 slope same, coefficient different, so parallel

OpenStudy (anonymous):

i meant coenciding, my bad, but your right